# Shield Synthesis for LTL Modulo Theories

### v1

@@ 1 Introduction
Recently, DNN-based agents trained using Deep Reinforcement Learning (DRL) have been shown to successfully control reactive systems of high complexity (e.g., (?)) , such as robotic platforms. However, despite their success, DRL controllers still suffer from various safety issues; e.g., small perturbations to their inputs, resulting either from noise or from a malicious adversary, can cause even state-of-the-art agents to react unexpectedly (e.g., (?)) . This issue raises severe concerns regarding the deployment of DRL-based agents in safety-critical reactive systems.
@@
In order to cope with these DNN reliability concerns, the formal methods community has recently put forth various tools and techniques that rigorously ensure the safe behaviour of DNNs (e.g., (?)) , and specifically, of DRL-controlled reactive systems (e.g., (?)) . One of the main approaches that is gaining popularity, is shielding (??), i.e., the incorporation of an external component (a “shield”) that forces an agent to behave safely according to a given specification. This specification $\varphi$ is usually expressed as a propositional formula, in which the atomic propositions represent the inputs ($I$) and outputs ($O$) of the system, controlled by the DNN in question. Once $\varphi$ is available, shielding seeks to guarantee that all behaviors of the given system $D$ satisfy $\varphi$ through the means of a shield $S$: whenever the system encounters an input $I$ that triggers an erroneous output (i.e., $O:D(I)$ for which $\varphi(I,O)$ does not hold), $S$ corrects $O$ and replaces it with another action $O^{\prime}$, to ensure that $\varphi$(I, O’) does hold. Thus, the combined system $D\cdot S$ never violates $\varphi$. Shields are appealing for multiple reasons: they do not require “white box” access to $D$, a single shield $S$ can be used for multiple variants of $D$, it is usually computationally cheaper than static methods like DNN verification etc. Moreover, shields are intuitive for practitioners, since they are synthesized based on the required $\varphi$. However, despite significant progress, modern shielding methods still suffer from a main setback: they are only applicable to specifications in which the inputs/outputs are over Boolean atomic propositions. This allows users to encode only discrete specifications, typically in Linear Temporal Logic (LTL). Thus, as most real-world systems rely on rich data specifications, this precludes the use of shielding in various such domains, such as continuous input spaces.
@@
In this work, we address this gap and present a novel approach for shield synthesis that makes use of LTL modulo theories ($\textup{LTL}_{\mathcal{T}}$), where Boolean propositions are extended to literals from a (multi-sorted) first-order theory $\mathcal{T}$. Concretely, we leverage Boolean abstraction methods (?), which transform $\textup{LTL}_{\mathcal{T}}$ specifications into equi-realizable pure (Boolean) LTL specifications. We combine Boolean abstraction with reactive $\textup{LTL}_{\mathcal{T}}$ synthesis (?), extending the common LTL shielding theory into a $\textup{LTL}_{\mathcal{T}}$ shielding theory. Using $\textup{LTL}_{\mathcal{T}}$ shielding, we are able to construct shields for more expressive specifications. This, in turn, allows us to override unwanted actions in a (possibly infinite) domain of $\mathcal{T}$, and guarantee the safety of DNN-controlled systems in such complex scenarios. In summary, our contributions are: (1) developing two methods for shield synthesis over $\textup{LTL}_{\mathcal{T}}$ and presenting their proof of correctness; (2) an analysis of the impact of the Boolean abstractions in the precision of shields; (3) a formalization of how to construct optimal shields using objective functions; and (4) an empirical evaluation that shows the applicability of our techniques.
2 Preliminaries
@@
We start from LTL (??), which has the following syntax:
@@
Reactive LTL synthesis (??) is the task of producing a system that satisfies a given LTL specification $\varphi$, where atomic propositions in $\varphi$ are split into variables controlled by the environment (“input variables”) and by the system (“output variables”), denoted by $\overline{e}$ and $\overline{s}$, respectively. Synthesis corresponds to a game where, in each turn, the environment player produces values for the input propositions, and the system player responds with values of the output propositions. A play is an infinite sequence of turns, i.e., an infinite interaction of the system with the environment. A strategy for the system is a tuple $C_{\mathbb{B}}:\langle Q,q_{0},\delta,o\rangle$ where $Q$ is a finite set of states, $q_{0}\in Q$ is the inital state, $\delta:Q\times\textit{val}(\overline{e})\mathrel{\rightarrow}Q$ is the transition function and $o:Q\times\textit{val}(\overline{e})\mathrel{\rightarrow}\textit{val}(\overline{s})$ is the output function. $C_{\mathbb{B}}$ is said to be winning for the system if all the possible plays played according to the strategy satisfy the LTL formula $\varphi_{\mathbb{B}}$. In this paper we use “strategy” and “controller” interchangeably.
@@
We also introduce the notion of winning region (WR), which encompasses all possible winning moves for the system in safety formulae. A winning region $\textit{WR}:\langle Q,I,T\rangle$ is a tuple where $Q$ is a finite set of states, $I\subseteq Q$ is a set of initial states and $T:Q\times\textit{val}(\overline{e})\mathrel{\rightarrow}{}2^{(Q\times\textit{val}(\overline{s}))}$ is the transition relation, which provides for a given state $q$ and input $v_{\overline{e}}$, all the possible pairs of legal successor and output $(q^{\prime},v_{\overline{s}})$. For a safety specification, every winning strategy is “included” into the WR (i.e., there is an embedding map). LTL realizability is the decision problem of whether there is a winning strategy for the system (i.e., check whether $\text{WR}\neq\emptyset$), while LTL synthesis is the computational problem of producing one.
@@
However, in $\textup{LTL}_{\mathcal{T}}$ synthesis (??), the specification is expressed in a richer logic where propositions are replaced by literals from some $\mathcal{T}$. In $\textup{LTL}_{\mathcal{T}}$ the (first order) variables in specification $\varphi_{\mathcal{T}}$ are still split into those controlled by the environment ($\overline{x}$), and those controlled by the system ($\overline{y}$), where $\overline{x}\cap\overline{y}=\emptyset$. We use $\varphi_{\mathcal{T}}(\overline{x},\overline{y})$ to empasize that $\overline{x}\cup\overline{y}$ are all the variables occurring in $\varphi_{\mathcal{T}}$. The alphabet is now $\Sigma_{\mathcal{T}}=\textit{val}(\overline{x}\cup\overline{y})$ (note that now valuations map a variable $x$ to $\mathbb{D}(\mathcal{T})$). We denote by $t[\overline{x}\leftarrow v_{\overline{x}}]$, the substitution in $t$ of variables $\overline{x}$ by values $v_{\overline{x}}$ (similarly for $t[\overline{y}\leftarrow v_{\overline{y}}]$), and also $t[\overline{e}\leftarrow v_{\overline{e}}]$ and $t[\overline{s}\leftarrow v_{\overline{s}}]$ for Boolean variables (propositions). A trace $\pi$ is an infinite sequence of valuations in $\mathbb{D}(\mathcal{T})$, which induces an infinite sequence of Boolean values of the literals occurring in $\varphi_{\mathcal{T}}$ and, in turn, an evaluation of $\varphi_{\mathcal{T}}$ using the semantics of the temporal operators. For example, given $\psi=\LTLsquare(y<x)$ the trace $\pi$$\{(x:2,y:6),(x:15,y:27)\ldots\}$ induces $\{(\textit{false}),(\textit{true})\ldots\}$. We use $\pi_{x}$ to denote the projection of $\pi$ to the values of only $x$ (resp. $\pi_{y}$ for $y$). A strategy or controller for the system in $\mathcal{T}$ is now a tuple $C_{\mathcal{T}}:\langle Q,q_{0},\delta,o\rangle$ where $Q$ and $q_{0}\in Q$ are as before, and $\delta:Q\times\textit{val}(\overline{x})\mathrel{\rightarrow}Q$ and $o:Q\times\textit{val}(\overline{x})\mathrel{\rightarrow}\textit{val}(\overline{y})$.
@@
Boolean abstraction (?) transforms an $\textup{LTL}_{\mathcal{T}}$ specification $\varphi_{\mathcal{T}}$ into an LTL specification $\varphi_{\mathbb{B}}$ in the same temporal fragment (e.g., safety to safety) that preserves realizability, i.e., $\varphi_{\mathcal{T}}$ and $\varphi_{\mathbb{B}}$ are equi-realizable. Then, $\varphi_{\mathbb{B}}$ can be fed into an off-the-shelf synthesis engine, which generates a controller or a WR for realizable instances. Boolean abstraction transforms $\varphi_{\mathcal{T}}$, which contains literals $l_{i}$, into $\varphi_{\mathbb{B}}=\varphi_{\mathcal{T}}[l_{i}\leftarrow s_{i}]\wedge\varphi^{\textit{extra}}$, where $\overline{s}=\{s_{i}|\text{for each }l_{i}\}$ is a set of fresh atomic propositions controlled by the system—such that $s_{i}$ replaces $l_{i}$—and where $\varphi^{\textit{extra}}$ is an additional sub-formula that captures the dependencies between the $\overline{s}$ variables. The formula $\varphi^{\textit{extra}}$ also includes additional variables $\overline{e}$ (controlled by the environment) that encode that the environment can leave the system with the power to choose certain valuations of the variables $\overline{s}$.
@@
We often represent a valuation $v_{\overline{s}}$ of the Boolean variables $\overline{s}$ (which map each variable in $\overline{s}$ to true or false) as a choice $c$ (an element of $2^{\overline{s}}$), where $s_{i}\in c$ means that $v_{\overline{s}}(s_{i})=\textit{true}$. The characteristic formula $f_{c}(\overline{x},\overline{y})$ of a choice $c$ is $f_{c}=\bigwedge_{s_{i}\in c}l_{i}\mathrel{\wedge}\bigwedge_{s_{i}\notin c}\neg l_{i}.$ We use $\mathcal{C}$ for the set of choices, i.e., sets of sets of $\overline{s}$. A reaction $r\subset\mathcal{C}$ is a set of choices, which characterizes the possible responses of the system as the result to a move of the environment. The characteristic formula $f_{r}(\overline{x})$ of a reaction $r$ is: $(\bigwedge_{c\in r}\exists\overline{y}.f_{c})\mathrel{\wedge}(\bigwedge_{c\notin r}\forall\overline{y}\neg f_{c})$. We say that $r$ is a valid reaction whenever $\exists\overline{x}.f_{r}(\overline{x})$ is valid. Intuitively, $f_{r}$ states that for some $v_{\overline{x}}$ by the environment, the system can respond with $v_{\overline{y}}$ making the literals in some choice $c\in r$ but cannot respond with $v_{\overline{y}}$ making the literals in choices $c\notin r$. The set VR of valid reactions partitions precisely the moves of the environment in terms of the reaction power left to the system. For each valid reaction $r$ there is a fresh environment variable $e\in\overline{e}$ used in $\varphi^{\textit{extra}}$ to capture the move of the environment that chooses reaction $r$. The formula $f_{r}(\overline{x})[\overline{x}\leftarrow v_{\overline{x}}]$ is true if a given valuation $v_{\overline{x}}$ of $\overline{x}$ is one move of the environment characterized by $r$. The formula $\varphi^{\textit{extra}}$ also constraints the environment such that exactly one of the variables in $\overline{e}$ is true, which forces that the environment chooses precisely one valid reaction $r$ when the variable $e$ that corresponds to $r$ is true.
@@
Let $\varphi_{\mathcal{T}}=\square(R_{0}\wedge R_{1})$ where:
@@
where $f_{c_{1}}=(s_{0}\wedge s_{1}\wedge\neg s_{2})$, $f_{c_{2}}=(s_{0}\wedge\neg s_{1}\wedge s_{2})$, $f_{c_{3}}=(s_{0}\wedge\neg s_{1}\wedge\neg s_{2})$, $f_{c_{4}}=(\neg s_{0}\wedge s_{1}\wedge s_{2})$, $f_{c_{5}}=(s_{0}\wedge s_{1}\wedge\neg s_{2})$ and $f_{c_{6}}=(\neg s_{0}\wedge\neg s_{1}\wedge s_{2})$ and where $c_{0}=\{s_{0},s_{1},s_{2}\}$, $c_{1}=\{s_{0},s_{1}\}$, $c_{2}=\{s_{0},s_{2}\}$, $c_{3}=\{s_{0}\}$, $c_{4}=\{s_{1},s_{2}\}$, $c_{5}=\{s_{1}\}$, $c_{6}=\{s_{1}\}$ and $c_{7}=\emptyset$. Also, $\overline{e}=\{e_{0},e_{0^{+}},e_{1}\}$ belong to the environment and represent $(x<10)$, $(x<9)$  [^1] We address implications of $(x<9)$ being subsumed by $(x<10)$ in Subsec. 3.2. and $(x\geq 10)$, respectively. $\varphi^{\textit{legal}}:(e_{0}\mathrel{\wedge}\neg e_{1}\mathrel{\wedge}\neg e_{2})\mathrel{\vee}(\neg e_{0}\mathrel{\wedge}e_{1}\mathrel{\wedge}\neg e_{2})\mathrel{\vee}(\neg e_{0}\mathrel{\wedge}\neg e_{1}\mathrel{\wedge}e_{2})$ encodes that $\overline{e}$ characterizes a partition of the (infinite) input valuations of the environment and that only one of the $\overline{e}$ is true in every move; e.g., the valuation $v_{\overline{e}}:\langle e_{0}:\textit{false},e_{0}^{+}:\textit{true},e_{1}:\textit{false}\rangle$ of $\overline{e}$ corresponds to the choice of the environment where only $e_{0}^{+}$ is true (and we use $v_{\overline{e}}:e_{0}^{+}$ for a shorter notation). Sub-formulae such as ($(\neg s_{0}\wedge s_{1}\wedge s_{2})$ represent the choices of the system (in this case, $c=\{s_{1},s_{2}\}$), that is, given a decision of the environment (a valuation of $\overline{e}$ that makes exactly one variable $e\in\overline{e}$ true), the system can react with one of the choices $c$ in the disjunction implied by $e$. Note that $f_{c}=\neg(x<2)\wedge(y>1)\wedge(y\leq x)$. Also, note that $c$ can be represented as $v_{\overline{s}}:\langle s_{0}:\textit{false},s_{1}:\textit{true},s_{2}:\textit{true}\rangle$.
3 Shielding with LTL Modulo Theories
@@ 3.1 $\textup{LTL}_{\mathcal{T}}$ Shield Construction
    Boolean abstraction (?) transforms $\varphi_{\mathcal{T}}$ into an equi-realizable LTL $\varphi_{\mathbb{B}}$.
@@
    A Boolean controller $C_{\mathbb{B}}$ is synthesized from $\varphi_{\mathbb{B}}$ (using e.g., (?)). $C_{\mathbb{B}}$ receives Boolean inputs $v_{\overline{e}}$ and provides Boolean outputs $v_{\overline{s}}$.
@@
We now elaborate on the last steps. To construct $C_{\mathcal{T}}$ from $C_{\mathbb{B}}$ we use (??) , where $C_{\mathcal{T}}$ is composed of three sub-components:
@@
  • A provider, which computes an output $v_{\overline{y}}^{\prime}$ from $f_{c}$ and $r$ such that $v_{\overline{y}}^{\prime}$ is a model [^2] In this paper we implement the calculation of $v_{\overline{y}}^{\prime}$ using SMT solving (concretely, Z3 (?)). An alternative is to statically use functional synthesis (such as AEval (?) for $\forall^{*}\exists^{*}$ formulae), which we study in detail in a companion paper and which does not allow optimization using soft constraints in Subsec. 3.3. of the formula $\exists\overline{y}^{\prime}.\big{(}f_{r}(\overline{x})\mathrel{\rightarrow}f_{c}(\overline{y}^{\prime},\overline{x})\big{)}[\overline{x}\leftarrow v_{\overline{x}}]$.
The composition of the partitioner, $C_{\mathbb{B}}$ and the provider implements $C_{\mathcal{T}}$ (see (??)). For each of these components $g$, let us denote with $g(t_{0})=t_{1}$ the application of $g$ with input $t_{0}$. Then, the following holds:
@@
In other words, given $\varphi_{\mathcal{T}}$, if $\mathcal{T}$ is decidable in the $\exists^{*}\forall^{*}$-fragment (a condition for the Boolean abstraction (?)), then, leveraging $C_{\mathcal{T}}$, a shield $S_{\mathcal{T}}$ conforming to $\varphi_{\mathcal{T}}$ can be computed for every external $D$ whose input-output domain is $\mathcal{T}$.
@@
The winning region (WR) (see Sec. 2) is the most general characterization of the winning states and moves for a safety LTL specification. This second construction can be applied to any subset of WR as long as every state has at least one successor for every input (for example, using the combination of a collection of controllers). The process of computing $S_{\mathcal{T}}$ based on a given winning region $W:\langle Q,I,T\rangle$ is as follows (see Fig. 1(b)), which maintains at each instant a set of current states $Q_{\textit{now}}\subseteq Q$ of $W$, starting from the initial set of states $I$.
@@
    We compute the winning region $W$ from $\varphi_{\mathbb{B}}$ (e.g., using (?)).

### v2

@@ Introduction
Recently, DNN-based agents trained using Deep Reinforcement Learning (DRL) have been shown to successfully control reactive systems of high complexity (e.g., (Marchesini and Farinelli 2020)) , such as robotic platforms. However, despite their success, DRL controllers still suffer from various safety issues; e.g., small perturbations to their inputs, resulting either from noise or from a malicious adversary, can cause even state-of-the-art agents to react unexpectedly (e.g., (Goodfellow, Shlens, and Szegedy 2014)) . This issue raises severe concerns regarding the deployment of DRL-based agents in safety-critical reactive systems.
@@
In order to cope with these DNN reliability concerns, the formal methods community has recently put forth various tools and techniques that rigorously ensure the safe behaviour of DNNs (e.g., (Katz et al. 2017)) , and specifically, of DRL-controlled reactive systems (e.g., (Bassan et al. 2023)) . One of the main approaches that is gaining popularity, is shielding (Bloem et al. 2015; Alshiekh et al. 2018), i.e., the incorporation of an external component (a “shield”) that forces an agent to behave safely according to a given specification. This specification $\varphi$ is usually expressed as a propositional formula, in which the atomic propositions represent the inputs ($I$) and outputs ($O$) of the system, controlled by the DNN in question. Once $\varphi$ is available, shielding seeks to guarantee that all behaviors of the given system $D$ satisfy $\varphi$ through the means of a shield $S$: whenever the system encounters an input $I$ that triggers an erroneous output (i.e., $O:D(I)$ for which $\varphi(I,O)$ does not hold), $S$ corrects $O$ and replaces it with another action $O^{\prime}$, to ensure that $\varphi$(I, O’) does hold. Thus, the combined system $D\cdot S$ never violates $\varphi$. Shields are appealing for multiple reasons: they do not require “white box” access to $D$, a single shield $S$ can be used for multiple variants of $D$, it is usually computationally cheaper than static methods like DNN verification etc. Moreover, shields are intuitive for practitioners, since they are synthesized based on the required $\varphi$. However, despite significant progress, modern shielding methods still suffer from a main setback: they are only applicable to specifications in which the inputs/outputs are over Boolean atomic propositions. This allows users to encode only discrete specifications, typically in Linear Temporal Logic (LTL). Thus, as most real-world systems rely on rich data specifications, this precludes the use of shielding in various such domains, such as continuous input spaces.
@@
In this work, we address this gap and present a novel approach for shield synthesis that makes use of LTL modulo theories ($\textup{LTL}_{\mathcal{T}}$), where Boolean propositions are extended to literals from a (multi-sorted) first-order theory $\mathcal{T}$. Concretely, we leverage Boolean abstraction methods (Rodriguez and Sánchez 2023), which transform $\textup{LTL}_{\mathcal{T}}$ specifications into equi-realizable pure (Boolean) LTL specifications. We combine Boolean abstraction with reactive $\textup{LTL}_{\mathcal{T}}$ synthesis (Rodriguez and Sánchez 2024), extending the common LTL shielding theory into a $\textup{LTL}_{\mathcal{T}}$ shielding theory. Using $\textup{LTL}_{\mathcal{T}}$ shielding, we are able to construct shields for more expressive specifications. This, in turn, allows us to override unwanted actions in a (possibly infinite) domain of $\mathcal{T}$, and guarantee the safety of DNN-controlled systems in such complex scenarios. In summary, our contributions are: (1) developing two methods for shield synthesis over $\textup{LTL}_{\mathcal{T}}$ and presenting their proof of correctness; (2) an analysis of the impact of the Boolean abstractions in the precision of shields; (3) a formalization of how to construct optimal shields using objective functions; and (4) an empirical evaluation that shows the applicability of our techniques.
Preliminaries
@@
We start from LTL (Pnueli 1977; Manna and Pnueli 1995), which has the following syntax:
@@
Reactive LTL synthesis (Thomas 2008; Piterman, Pnueli, and Sa’ar 2006) is the task of producing a system that satisfies a given LTL specification $\varphi$, where atomic propositions in $\varphi$ are split into variables controlled by the environment (“input variables”) and by the system (“output variables”), denoted by $\overline{e}$ and $\overline{s}$, respectively. Synthesis corresponds to a game where, in each turn, the environment player produces values for the input propositions, and the system player responds with values of the output propositions. A play is an infinite sequence of turns, i.e., an infinite interaction of the system with the environment. A strategy for the system is a tuple $C_{\mathbb{B}}:\langle Q,q_{0},\delta,o\rangle$ where $Q$ is a finite set of states, $q_{0}\in Q$ is the inital state, $\delta:Q\times\textit{val}(\overline{e})\mathrel{\rightarrow}Q$ is the transition function and $o:Q\times\textit{val}(\overline{e})\mathrel{\rightarrow}\textit{val}(\overline{s})$ is the output function. $C_{\mathbb{B}}$ is said to be winning for the system if all the possible plays played according to the strategy satisfy the LTL formula $\varphi_{\mathbb{B}}$. In this paper we use “strategy” and “controller” interchangeably.
@@
We also introduce the notion of winning region (WR), which encompasses all possible winning moves for the system in safety formulae. A winning region $\textit{WR}:\langle Q,I,T\rangle$ is a tuple where $Q$ is a finite set of states, $I\subseteq Q$ is a set of initial states and $T:Q\times\textit{val}(\overline{e})\mathrel{\rightarrow}{}2^{(Q\times\textit{val}(\overline{s}))}$ is the transition relation, which provides for a given state $q$ and input $v_{\overline{e}}$, all the possible pairs of legal successor and output $(q^{\prime},v_{\overline{s}})$. For a safety specification, every winning strategy is “included” into the WR (i.e., there is an embedding map). LTL realizability is the decision problem of whether there is a winning strategy for the system (i.e., check if $\text{WR}\neq\emptyset$), while LTL synthesis is the computational problem of producing one.
@@
However, in $\textup{LTL}_{\mathcal{T}}$ synthesis (Rodríguez and Sánchez 2024; Rodriguez and Sánchez 2024), the specification is expressed in a richer logic where propositions are replaced by literals from some $\mathcal{T}$. In $\textup{LTL}_{\mathcal{T}}$ the (first order) variables in specification $\varphi_{\mathcal{T}}$ are still split into those controlled by the environment ($\overline{x}$), and those controlled by the system ($\overline{y}$), where $\overline{x}\cap\overline{y}=\emptyset$. We use $\varphi_{\mathcal{T}}(\overline{x},\overline{y})$ to empasize that $\overline{x}\cup\overline{y}$ are all the variables occurring in $\varphi_{\mathcal{T}}$. The alphabet is now $\Sigma_{\mathcal{T}}=\textit{val}(\overline{x}\cup\overline{y})$ (note that now valuations map a variable $x$ to $\mathbb{D}(\mathcal{T})$). We denote by $t[\overline{x}\leftarrow v_{\overline{x}}]$, the substitution in $t$ of variables $\overline{x}$ by values $v_{\overline{x}}$ (similarly for $t[\overline{y}\leftarrow v_{\overline{y}}]$), and also $t[\overline{e}\leftarrow v_{\overline{e}}]$ and $t[\overline{s}\leftarrow v_{\overline{s}}]$ for Boolean variables (propositions). A trace $\pi$ is an infinite sequence of valuations in $\mathbb{D}(\mathcal{T})$, which induces an infinite sequence of Boolean values of the literals occurring in $\varphi_{\mathcal{T}}$ and, in turn, an evaluation of $\varphi_{\mathcal{T}}$ using the semantics of the temporal operators. For example, given $\psi=\LTLsquare(y<x)$ the trace $\pi$$\{(x:2,y:6),(x:15,y:27)\ldots\}$ induces $\{(\textit{false}),(\textit{true})\ldots\}$. We use $\pi_{x}$ to denote the projection of $\pi$ to the values of only $x$ (resp. $\pi_{y}$ for $y$). A strategy or controller for the system in $\mathcal{T}$ is now a tuple $C_{\mathcal{T}}:\langle Q,q_{0},\delta,o\rangle$ where $Q$ and $q_{0}\in Q$ are as before, and $\delta:Q\times\textit{val}(\overline{x})\mathrel{\rightarrow}Q$ and $o:Q\times\textit{val}(\overline{x})\mathrel{\rightarrow}\textit{val}(\overline{y})$.
@@
Boolean abstraction (Rodriguez and Sánchez 2023) transforms an $\textup{LTL}_{\mathcal{T}}$ specification $\varphi_{\mathcal{T}}$ into an LTL specification $\varphi_{\mathbb{B}}$ in the same temporal fragment (e.g., safety to safety) that preserves realizability, i.e., $\varphi_{\mathcal{T}}$ and $\varphi_{\mathbb{B}}$ are equi-realizable. Then, $\varphi_{\mathbb{B}}$ can be fed into an off-the-shelf synthesis engine, which generates a controller or a WR for realizable instances. Boolean abstraction transforms $\varphi_{\mathcal{T}}$, which contains literals $l_{i}$, into $\varphi_{\mathbb{B}}=\varphi_{\mathcal{T}}[l_{i}\leftarrow s_{i}]\wedge\varphi^{\textit{extra}}$, where $\overline{s}=\{s_{i}|\text{for each }l_{i}\}$ is a set of fresh atomic propositions controlled by the system—such that $s_{i}$ replaces $l_{i}$—and where $\varphi^{\textit{extra}}$ is an additional sub-formula that captures the dependencies between the $\overline{s}$ variables. The formula $\varphi^{\textit{extra}}$ also includes additional environment variables $\overline{e}$ that encode that the environment can leave the system with the power to choose certain valuations of the variables $\overline{s}$.
@@
We often represent a valuation $v_{\overline{s}}$ of the Boolean variables $\overline{s}$ (which map each variable in $\overline{s}$ to true or false) as a choice $c$ (an element of $2^{\overline{s}}$), where $s_{i}\in c$ means that $v_{\overline{s}}(s_{i})=\textit{true}$. The characteristic formula $f_{c}(\overline{x},\overline{y})$ of a choice $c$ is $f_{c}=\bigwedge_{s_{i}\in c}l_{i}\mathrel{\wedge}\bigwedge_{s_{i}\notin c}\neg l_{i}.$ We use $\mathcal{C}$ for the set of choices, i.e., sets of sets of $\overline{s}$. A reaction $r\subset\mathcal{C}$ is a set of choices, which characterizes the possible responses of the system as the result to a move of the environment. The characteristic formula $f_{r}(\overline{x})$ of a reaction $r$ is: $(\bigwedge_{c\in r}\exists\overline{y}.f_{c})\mathrel{\wedge}(\bigwedge_{c\notin r}\forall\overline{y}\neg f_{c})$. We say that $r$ is a valid reaction whenever $\exists\overline{x}.f_{r}(\overline{x})$ is valid. Intuitively, $f_{r}$ states that for some $v_{\overline{x}}$ by the environment, the system can respond with $v_{\overline{y}}$ making the literals in some choice $c\in r$ but cannot respond with $v_{\overline{y}}$ making the literals in choices $c\notin r$. The set VR of valid reactions partitions precisely the moves of the environment in terms of the reaction power left to the system. For each valid reaction $r$ there is a fresh environment variable $e\in\overline{e}$ used in $\varphi^{\textit{extra}}$ to capture the move of the environment that chooses reaction $r$. The formula $f_{r}(\overline{x})[\overline{x}\leftarrow v_{\overline{x}}]$ is true if a given valuation $v_{\overline{x}}$ of $\overline{x}$ is one move of the environment characterized by $r$. The formula $\varphi^{\textit{extra}}$ restricts the environment such that exactly one of the variables in $\overline{e}$ is true, which forces that the environment chooses precisely one valid reaction $r$ when the variable $e$ that corresponds to $r$ is true.
@@
In this paper, we address shields for general safety $\textup{LTL}_{\mathcal{T}}$, but for the sake of simplicity in the presentation, we consider a simpler example. Let $\varphi_{\mathcal{T}}=\square(R_{0}\wedge R_{1})$ where:
@@
where $f_{c_{1}}=(s_{0}\wedge s_{1}\wedge\neg s_{2})$, $f_{c_{2}}=(s_{0}\wedge\neg s_{1}\wedge s_{2})$, $f_{c_{3}}=(s_{0}\wedge\neg s_{1}\wedge\neg s_{2})$, $f_{c_{4}}=(\neg s_{0}\wedge s_{1}\wedge s_{2})$, $f_{c_{5}}=(s_{0}\wedge s_{1}\wedge\neg s_{2})$ and $f_{c_{6}}=(\neg s_{0}\wedge\neg s_{1}\wedge s_{2})$ and where $c_{0}=\{s_{0},s_{1},s_{2}\}$, $c_{1}=\{s_{0},s_{1}\}$, $c_{2}=\{s_{0},s_{2}\}$, $c_{3}=\{s_{0}\}$, $c_{4}=\{s_{1},s_{2}\}$, $c_{5}=\{s_{1}\}$, $c_{6}=\{s_{1}\}$ and $c_{7}=\emptyset$. Also, $\overline{e}=\{e_{0},e_{0^{+}},e_{1}\}$ belong to the environment and represent $(x<10)$, $(x<9)$ and $(x\geq 10)$, respectively. $\varphi^{\textit{legal}}:(e_{0}\mathrel{\wedge}\neg e_{1}\mathrel{\wedge}\neg e_{2})\mathrel{\vee}(\neg e_{0}\mathrel{\wedge}e_{1}\mathrel{\wedge}\neg e_{2})\mathrel{\vee}(\neg e_{0}\mathrel{\wedge}\neg e_{1}\mathrel{\wedge}e_{2})$ encodes that $\overline{e}$ characterizes a partition of the (infinite) input valuations of the environment and that only one of the $\overline{e}$ is true in every move; e.g., the valuation $v_{\overline{e}}:\langle e_{0}:\textit{false},e_{0}^{+}:\textit{true},e_{1}:\textit{false}\rangle$ of $\overline{e}$ corresponds to the choice of the environment where only $e_{0}^{+}$ is true (and we use $v_{\overline{e}}:e_{0}^{+}$ for a shorter notation). Sub-formulae such as ($(\neg s_{0}\wedge s_{1}\wedge s_{2})$ represent the choices of the system (in this case, $c=\{s_{1},s_{2}\}$), that is, given a decision of the environment (a valuation of $\overline{e}$ that makes exactly one variable $e\in\overline{e}$ true), the system can react with one of the choices $c$ in the disjunction implied by $e$. Note that $f_{c}=\neg(x<2)\wedge(y>1)\wedge(y\leq x)$. Also, note that $c$ can be represented as $v_{\overline{s}}:\langle s_{0}:\textit{false},s_{1}:\textit{true},s_{2}:\textit{true}\rangle$.
$\textup{LTL}_{\mathcal{T}}$ Shield Computation
Problem overview.
@@
    Boolean abstraction (Rodriguez and Sánchez 2023) transforms $\varphi_{\mathcal{T}}$ into an equi-realizable LTL $\varphi_{\mathbb{B}}$.
@@
    A Boolean controller $C_{\mathbb{B}}$ is synthesized from $\varphi_{\mathbb{B}}$ (using e.g., (Meyer, Sickert, and Luttenberger 2018)). $C_{\mathbb{B}}$ receives Boolean inputs $v_{\overline{e}}$ and provides Boolean outputs $v_{\overline{s}}$.
@@
We now elaborate on the last steps. To construct $C_{\mathcal{T}}$ from $C_{\mathbb{B}}$ we use (Rodríguez and Sánchez 2024) , where $C_{\mathcal{T}}$ is composed of three sub-components:
@@
  • A provider, which computes an output $v_{\overline{y}}^{\prime}$ from $f_{c}$ and $r$ such that $v_{\overline{y}}^{\prime}$ is a model See the extended version of this paper (Rodriguez et al. 2024) for a for additional details. of the formula $\exists\overline{y}^{\prime}.\big{(}f_{r}(\overline{x})\mathrel{\rightarrow}f_{c}(\overline{y}^{\prime},\overline{x})\big{)}[\overline{x}\leftarrow v_{\overline{x}}]$.
The composition of the partitioner, $C_{\mathbb{B}}$ and the provider implements $C_{\mathcal{T}}$ (see (Rodríguez and Sánchez 2024; Rodriguez and Sánchez 2024)). For each of these components $g$, let us denote with $g(t_{0})=t_{1}$ the application of $g$ with input $t_{0}$. Then, the following holds:
@@
In other words, given $\varphi_{\mathcal{T}}$, if $\mathcal{T}$ is decidable in the $\exists^{*}\forall^{*}$-fragment (a condition for the Boolean abstraction (Rodriguez and Sánchez 2023)), then, leveraging $C_{\mathcal{T}}$, a shield $S_{\mathcal{T}}$ conforming to $\varphi_{\mathcal{T}}$ can be computed for every external $D$ whose input-output domain is $\mathcal{T}$.
@@
The winning region (WR) is the most general characterization of the winning states and moves for a safety LTL specification. This second construction can be applied to any subset of WR as long as every state has at least one successor for every input (for example, using the combination of a collection of controllers). The process of computing $S_{\mathcal{T}}$ based on a given winning region $W:\langle Q,I,T\rangle$ is as follows (see Fig. 1(b)) and maintains at each instant a set of current states $Q_{\textit{now}}\subseteq Q$ of $W$, starting from the initial set of states $I$.
@@
    We compute the winning region $W$ from $\varphi_{\mathbb{B}}$ (e.g., using (Brenguier et al. 2014)).